# Imagine-then-Plan: Agent Learning from Adaptive Lookahead with World Models

### v1

@@ Abstract
Recent advances in world models have shown promise for modeling future dynamics of environmental states, enabling agents to reason and act without accessing real environments. Current methods mainly perform single-step or fixed-horizon rollouts, leaving their potential for complex task planning under-exploited. We propose Imagine-then-Plan (ITP), a unified framework for agent learning via lookahead imagination, where an agent’s policy model interacts with the learned world model, yielding multi-step “imagined” trajectories. Since the imagination horizon may vary by tasks and stages, we introduce a novel adaptive lookahead mechanism by trading off the ultimate goal and task progress. The resulting imagined trajectories provide rich signals about future consequences, such as achieved progress and potential conflicts, which are fused with current observations, formulating a partially observable and imaginable Markov decision process to guide policy learning. We instantiate ITP with both training-free and reinforcement-trained variants. Extensive experiments across representative agent benchmarks demonstrate that ITP significantly outperforms competitive baselines. Further analyses validate that our adaptive lookahead largely enhances agents’ reasoning capability, providing valuable insights into addressing broader, complex tasks. Our code and data will be released.
@@ 1 Introduction
The emergence of Large Language Models (LLMs) has sparked a paradigm shift in autonomous agents, enabling them to reason and interact across a wide range of digital and physical environments (Li et al., 2023; Wang et al., 2024a; Fung et al., 2025). LLM-based agents primarily leverage immediate observations and historical interaction traces to facilitate decision-making (Yao et al., 2023; Shinn et al., 2023). Despite the impressive performance achieved, most agents remain constrained by shallow grounding, a state where they perceive the environment but lack a deep, causal understanding of how their current actions will ultimately reshape the environment. Without the ability to project into the future, agents are prone to catastrophic failures, discovering erroneous actions or state conflicts only after they have been irreversibly executed.
@@
To bridge this grounding gap, world models (LeCun, 2022; Hafner et al., 2023) that focus on future state modeling have emerged as a promising solution, enabling agents to simulate environmental dynamics and “rehearse” actions in a mental sandbox (Wang et al., 2025d). Contemporary methods typically employ world models for single-step verification or with fixed-horizon rollouts Hao et al. (2023); Qiao et al. (2024); Zhang et al. (2025). Such rigid strategies are often suboptimal, as they fail to capture long-term dependencies in complex tasks (e.g., household embodied tasks). Furthermore, they are prone to suffer from high computational costs with unnecessary deep rollouts for trivial actions. A truly intelligent agent should be deliberative, allocating deep foresight adaptively to resolve potential state conflicts and account for long-term dependencies in pivotal, high-stakes decisions.
@@ 2 Preliminaries
Figure 2: Overview of the proposed Imageine-then-Plan (ITP) framework. It consists of two variants: (a) $\texttt{ITP}_{\text{I}}$, which is training-free and enables LLM agents to learn from the imagination at inference time. (b) $\texttt{ITP}_{\text{R}}$, which leverages imagined futures to optimize the action policy more effectively and more efficiently.
@@ 3.1 World Model Training
As shown in Figure 2, we first fine-tune a base LLM on expert demonstrations $\mathcal{D}_{\mathrm{exp}}$ to obtain an initial agent policy $\pi_{\theta_{0}}$. This warm-up establishes basic capability to produce executable actions, serving as the foundation for the agent’s exploration. We ask the agent to perform rollouts in the environment, obtaining the rollout trajectories $\mathcal{D}_{\mathrm{roll}}$. As introduced in § 2, we learn a LLM-based world model $\mathcal{M}_{\phi}$ that approximates the environment dynamics $p_{\phi}(s{{}^{\prime}}|s,a)$, where $s$ and $a$ denote the current state and action, respectively. $s{{}^{\prime}}$ represents the next-step state. To ensure the world model is grounded and robust to out-of-distribution actions, we train it on a joint dataset $\mathcal{D_{\mathrm{WM}}}=\mathcal{D}_{\mathrm{exp}}\cup\mathcal{D}_{\mathrm{roll}}$. The world model $\mathcal{M}_{\phi}$ is optimized by minimizing the negative log-likelihood as follows:
@@ 3.3 Planning with Adaptive Lookahead
To achieve effective task planning, a key challenge is determining how far the agent should imagine. Short-horizon imagination may miss long-term dependencies, while excessive rollouts can amplify model errors and incur unnecessary computation. To resolve this, we aim to adaptively select the imagination horizon $K_{t}$ based on the estimated task progress against the ultimate goal. We instantiate ITP via two distinct variants to provide both flexibility and optimization: (i) $\texttt{ITP}_{\text{I}}$, which is an inference-time method that learns from the imagination, and (ii) $\texttt{ITP}_{\text{R}}$, which is a reinforcement-trained method that jointly optimizes the lookahead horizon selection and action policy.
@@ 3.3.1 In-Imagination Learning ($\texttt{ITP}_{\text{I}}$)
$\texttt{ITP}_{\text{I}}$ is a training-free variant designed for improving LLM agents during inference time. In this mode, both the agent policy and the world model remain frozen, and the agent relies on its inherent capabilities to perform deliberative reasoning. At each step $t$, the agent receives the current state $s_{t}$ and executes a three-stage “Imagine-then-Plan” procedure: 1) Adaptive horizon selection: The agent first assesses the task instruction and the current state $s_{t}$ to determine an appropriate imagination horizon $K_{t}\in\{0,1,\dots,K_{\max}\}$. This allows the agent to allocate deeper foresight to high-stakes decisions while maintaining computational efficiency for trivial actions. 2) World-model imagination: The agent invokes its world model to perform a “mental rehearsal”. By interacting with $\mathcal{M}_{\phi}$ for $K_{t}$ steps, the agent obtains a multi-step future trajectory $\hat{\tau}_{t}^{(K_{t})}$. 3) Reflective policy generation: Rather than directly taking the first action of the imagined sequence, the agent uses $\hat{\tau}_{t}^{(K_{t})}$ as implicit feedback for reflective self-refinement.
Specifically, the agent is prompted to self-reflect on the imagined trajectory to evaluate task progress, determine if the predicted states move closer to the ultimate task goal, and detect potential conflicts, bottlenecks, or catastrophic failures before they are irreversibly executed in the real environment. The agent is asked to perform self-refinement and then select the optimal next action $a_{t}$, given by:
@@
By grounding decisions in these imagined futures, $\texttt{ITP}_{\text{I}}$ transforms passive observation into proactive deliberation, significantly improving task success rates without additional training.
@@ 3.3.2 Reinforced Training ($\texttt{ITP}_{\text{R}}$)
Compared to $\texttt{ITP}_{\text{I}}$ that relies on inference-time reasoning, $\texttt{ITP}_{\text{R}}$ aims to explicitly learn when and how long to imagine. We augment the agent with a lightweight $K$-head predictor $P_{\theta}(K_{t}|s_{t})$, which is a linear layer built on top of a backbone LLM that predicts distributions over imagination horizons. The action policy and predictor are optimized jointly with the following three stages.
@@
where $\alpha$ and $\beta$ are hyperparameters balancing value estimation and exploration. Algorithm 1 shows the overview of $\texttt{ITP}_{\text{R}}$’s training process.

### v2

@@ Abstract
Recent advances in world models have shown promise for modeling future dynamics of environmental states, enabling agents to reason and act without accessing real environments. Current methods mainly perform single-step or fixed-horizon rollouts, leaving their potential for complex task planning under-exploited. We propose Imagine-then-Plan (ITP), a unified framework for agent learning via lookahead imagination, where an agent’s policy model interacts with the learned world model, yielding multi-step “imagined” trajectories. Since the imagination horizon may vary by tasks and stages, we introduce a novel adaptive lookahead mechanism by trading off the ultimate goal and task progress. The resulting imagined trajectories provide rich signals about future consequences, such as achieved progress and potential conflicts, which are fused with current observations, formulating a partially observable and imaginable Markov decision process to guide policy learning. We instantiate ITP with both training-free and reinforcement-trained variants. Extensive experiments across representative agent benchmarks demonstrate that ITP significantly outperforms competitive baselines. Further analyses validate that our adaptive lookahead largely enhances agents’ reasoning capability, providing valuable insights into addressing broader, complex tasks. Our code and data will be publicly available at: https://github.com/loyiv/ITP.
@@ 1 Introduction
The emergence of Large Language Models (LLMs) has sparked a paradigm shift in autonomous agents, enabling them to reason and interact across a wide range of digital and physical environments (Li et al., 2023; Wang et al., 2024a; Fung et al., 2025). LLM-based agents primarily leverage immediate observations and historical interaction traces to facilitate decision-making (Yao et al., 2023b; Shinn et al., 2023). Despite the impressive performance achieved, most agents remain constrained by shallow grounding, a state where they perceive the environment but lack a deep, causal understanding of how their current actions will ultimately reshape the environment. Without the ability to project into the future, agents are prone to catastrophic failures, discovering erroneous actions or state conflicts only after they have been irreversibly executed.
@@
To bridge this grounding gap, world models (LeCun, 2022; Hafner et al., 2023) that focus on future state modeling have emerged as a promising solution, enabling agents to simulate environmental dynamics and “rehearse” actions in a mental sandbox (Wang et al., 2025e). Contemporary methods typically employ world models for single-step verification or with fixed-horizon rollouts Hao et al. (2023); Qiao et al. (2024); Zhang et al. (2025). Such rigid strategies are often suboptimal, as they fail to capture long-term dependencies in complex tasks (e.g., household embodied tasks). Furthermore, they are prone to suffer from high computational costs with unnecessary deep rollouts for trivial actions. A truly intelligent agent should be deliberative, allocating deep foresight adaptively to resolve potential state conflicts and account for long-term dependencies in pivotal, high-stakes decisions.
@@ 2 Preliminaries
Figure 2: Overview of the proposed Imagine-then-Plan (ITP) framework. It consists of two variants: (a) $\texttt{ITP}_{\text{I}}$, which is training-free and enables LLM agents to learn from the imagination at inference time. (b) $\texttt{ITP}_{\text{R}}$, which leverages an imagined future to optimize the action policy more effectively and more efficiently.
@@ 3.1 World Model Training
As shown in Figure 2, we first fine-tune a base LLM on expert demonstrations $\mathcal{D}_{\mathrm{exp}}$ to obtain an initial agent policy $\pi_{\theta_{0}}$. This warm-up establishes basic capability to produce executable actions, serving as the foundation for the agent’s exploration. We ask the agent to perform rollouts in the environment, obtaining the rollout trajectories $\mathcal{D}_{\mathrm{roll}}$. As introduced in § 2, we learn a LLM-based world model $\mathcal{M}_{\phi}$ that approximates the environment dynamics $p_{\phi}(s{{}^{\prime}}|s,a)$, where $s$ and $a$ denote the current state and action, respectively. $s{{}^{\prime}}$ represents the next-step state. To ensure the world model is grounded and robust to out-of-distribution actions, we train it on a joint dataset $\mathcal{D_{\mathrm{WM}}}=\mathcal{D}_{\mathrm{exp}}\cup\mathcal{D}_{\mathrm{roll}}$.
The world model $\mathcal{M}_{\phi}$ is optimized by minimizing the negative log-likelihood as follows:
@@
We provide detailed training setup, dataset scale, and computational costs in Appendix C.1.
@@ 3.3 Planning with Adaptive Lookahead
For effective task planning, a key challenge is determining how far the agent should imagine. Short-horizon imagination may miss long-term dependencies, while excessive rollouts can amplify model errors and incur unnecessary computation. To resolve this, we aim to adaptively select the imagination horizon $K_{t}$ based on the estimated task progress against the ultimate goal. We instantiate ITP via two distinct variants to provide both flexibility and optimization: (i) $\texttt{ITP}_{\text{I}}$, which is an inference-time method that learns from the imagination, and (ii) $\texttt{ITP}_{\text{R}}$, which is a reinforcement-trained method that jointly optimizes the lookahead horizon selection and action policy.
@@ 3.3.1 In-Imagination Learning ($\texttt{ITP}_{\text{I}}$)
$\texttt{ITP}_{\text{I}}$ is a training-free variant that improves LLM agents at inference time. Both the policy and world model remain frozen, and the agent performs deliberative reasoning based on the current state $s_{t}$. At each step $t$, it follows a three-stage “Imagine-then-Plan” procedure: 1) Adaptive horizon selection: The agent analyzes the instruction and state $s_{t}$ to choose an imagination horizon $K_{t}\in\{0,1,\dots,K_{\max}\}$, assigning deeper foresight to critical decisions while avoiding unnecessary computation on trivial ones. 2) World-model imagination: The agent interacts with $\mathcal{M}_{\phi}$ for $K_{t}$ steps to obtain a future trajectory $\hat{\tau}_{t}^{(K_{t})}$. 3) Reflective policy generation: Instead of directly executing the first imagined action, the agent treats $\hat{\tau}_{t}^{(K_{t})}$ as implicit feedback for self-refinement.
Specifically, the agent reflects on the imagined trajectory to assess progress toward the goal and identify potential conflicts, bottlenecks, or catastrophic failures before acting in the real environment. It then refines its reflection and selects the next optimal action $a_{t}$ as follows:
@@
By grounding decisions in the imagined future, $\texttt{ITP}_{\text{I}}$ turns passive observation into proactive, deliberative learning, improving task success without additional training.
@@ 3.3.2 Reinforced Training ($\texttt{ITP}_{\text{R}}$)
$\texttt{ITP}_{\text{R}}$ aims to explicitly learn when and how long to imagine. We augment the agent with a lightweight $K$-head predictor $P_{\theta}(K_{t}|s_{t})$, which is a linear layer built on top of a backbone LLM that predicts distributions over imagination horizons. The action policy and predictor are optimized jointly with the following three stages.
@@
where $\alpha$ and $\beta$ are hyperparameters balancing value estimation and exploration. Algorithm 1 shows the pseudocode of ITPR’s training process.